OpenStudy (anonymous):

Alex stores fertilizer for his farm in a rectangular box. The length of the box is 6 feet, width is 3 inches, and height is 9 inches. Each bag of fertilizer contains 0.225 cubic foot of fertilizer. How many bags of fertilizer will fill the box completely? [1 foot = 12 inches] Numerical Answers Expected! Answer for Blank 1: i think its 5

OpenStudy (anonymous):

@imqwerty

OpenStudy (anonymous):

1 foot =12 inches 1/12 foot=1inch So 3inches=1/12 *3=1/4foot 9inchees=1/12 *9=3/4foot 6ft * 0.25ft * 0.75ft = 1.125 Then divide that by 0.225.... 1.125 / 0.225 = 5

imqwerty (imqwerty):

yes correct :)